OpenStudy (anonymous):

A film shop carrying black-and-white film and color film has $4000 in inventory. The profit on black-and-white film is 12%, and the profit on color film is 21%. If all the film is sold, and if the profit on color film is $150 less than the profit on black-and-white film, how much was invested in each type of film?

OpenStudy (anonymous):

Let bw and c be the amount of money invested in black-and-white and color film respectively. Solve the following equations for bw and c:\[\left\{\text{bw}+c=4000,\frac{12}{100} \text{bw}-\frac{21}{100}c=150\right\}\]\[\{\text{bw}=\text{$\$$3000},c=\text{$\$$1000}\} \]Not requested in the problem statement, the gross sales amount from the sale of the $4000 inventory would be $4570.

OpenStudy (anonymous):

Thank you!